A certain recipe calls for 8 and 3/4 cups of sugar. If the recipe is to be doubled, how much sugar should be used? (Enter your answer as a proper fraction or mixed number, not as an improper fraction.)
Answer by checkley77(12844)

2(8 3/4)=2*35/4=35/2 ans.